OpenStudy (anonymous):

Can someone help me work out 3/8 + 1/2 and write how they worked it out?

OpenStudy (anonymous):

3/8+1/2=3/8+4/8=7/8...here for the sum to be worked out the denominators should be equal.it can be done by taking the lcm..here the lcm is 8.so to make the denominator equal multiply the denominator of 1/2 by 4..now the denominators are same so we can add them.ie,3/8+4/8=7/8...this is the answer...hopes so..if its wrong pls correct me..:)